Earlier reports Convert Day of Year

The Hubble Space Telescope

Unlocking the Secrets of the Universe

Daily Report #6702

Period Covered:
07:00 pm November 18, 2014 - 06:59 pm November 19, 2014
(DOY 323/0000z - 323/2359z)

Flight Operations Summary

This report was generated by the HST Automated Report System @2014-11-20 17:02:00 GMT

Significant Spacecraft Anomalies: (The following are preliminary reports
of potential non-nominal performance that will be investigated.)




FGS GSAcq		10		10
FGS REAcq		7		7
OBAD with Maneuver	9		9

A tabletop briefing to update was held on Tuesday November 18 to review plans for the update of the 
TDRS element for use during the EFT-1 launch and authority to proceed was received.  The activities 
will proceed throughout the installation period as detailed in the briefing package. Ops Request 
19704-1 – “TDRS 150 Orbital Element Update” and GSCA #2538-0 – “TDRS Usage File Update 
for TDRS-150” detail the steps that will be taken for the installation. An additional TDRS 
tabletop Ops briefing will be held in December to detail the steps to be taken after EFT-1 launch. 

Ops Request 19704-1 to uplink TDRS-150 (F-5) orbital elements into HST486 slot A was successfully 
executed from 323/15:02 – 15:09z.  A 486 memory dump of table 224 (TDRS Elements table) was 
performed and the dump was verified by 486 FSW.  TDRS-150 elements were uplinked to provide 
additional scheduling resources for HST during the Exploration Flight Test -1 (EFT-1) mission 
scheduled to launch on December 4th at 1205z.  TDRS resources will be very limited during the EFT-1 
launch window and flight period, which could be up to a maximum of approximately 8.7 hours 
beginning at T-30 minutes

Ops Request 19705 was successfully completed approximately ~323/16:53 to dump the first 6 banks of 
NSSC-1 memory . This periodic memory dump will be used to compile a database of nominal NSSC-1 
states for Payload FSW record keeping and analysis.

Both SAS1 and ALT1 have been checked out as being back in primary status.

Program Principal Investigator Program Title
Science Observations Scheduled
13645 Xiaohui Fan, University of Arizona Galactic Environment of A Twenty-Billion Solar-Mass Black Hole at the End of Reionization
13657 Jeyhan Kartaltepe, National Optical Astronomy Observatory, AURA Probing the Most Luminous Galaxies in the Universe at the Peak of Galaxy Assembly
13679 Lorenz Roth, Royal Institute of Technology Europa's Water Vapor Plumes: Systematically Constraining their Abundance and Variability
13774 Sara Ellison, University of Victoria Feeding and feeback: The impact of AGN on the circumgalactic medium.
13776 Michael Gregg, University of California - Davis Completing The Next Generation Spectral Library
13793 Rebecca Bowler, Royal Observatory Edinburgh Unveiling the merger fraction, sizes and morphologies of the brightest z ~ 7 galaxies
13868 Dale Kocevski, Colby College Are Compton-Thick AGN the Missing Link Between Mergers and Black Hole Growth?
Calibration Observations Scheduled
13526 Steven Penton, Space Telescope Science Institute COS Imaging TA and Spectroscopic WCA-PSA/BOA offset verifications
13952 David Golimowski, Space Telescope Science Institute CCD Daily Monitor (Part 1)
13957 David Golimowski, Space Telescope Science Institute CCD Hot Pixel Annealing
13980 Joanna Taylor, Space Telescope Science Institute CCD Dark Monitor Part 1
13982 Joanna Taylor, Space Telescope Science Institute CCD Bias and Read Noise Monitor Part 1
14002 Matthew Bourque, Space Telescope Science Institute WFC3 UVIS Daily Monitor A